OpenStudy (anonymous):

find the missing dimension of the cone. round to nerest tenth...volume=1/18 pi ft^3

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

sorry it show the bottom is 2/3 completely across

OpenStudy (anonymous):

so 1/3 right?

OpenStudy (luigi0210):

The radius?

OpenStudy (anonymous):

yes. missing height

OpenStudy (anonymous):

and one more and I'll post after

OpenStudy (luigi0210):

Well let's work backwards.. http://prntscr.com/3a4oy2 \[\LARGE \frac{1}{18} \pi= \pi r^2 \frac{h}{3}\]

OpenStudy (luigi0210):

And you said 2/3 was all the way across?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

So yea, 1/3=r Plug that in too.. \[\LARGE \frac{1}{18} \pi= \pi (\frac{1}{3})^2 \frac{h}{3}\] Simpify it \[\LARGE \frac{1}{18} \pi= \frac{\pi h}{18}\]

OpenStudy (luigi0210):

Can you finish from here?

OpenStudy (anonymous):

umm can you help?

OpenStudy (anonymous):

not real good at math

OpenStudy (luigi0210):

We have \[\LARGE \frac{\pi}{18}= \frac{\pi h}{18}\] We just multiply by the reciprocal like so.. \[\LARGE \frac{\pi}{18}*\frac{18}{\pi} = \frac{\pi h}{18}*\frac{18}{\pi}\] That'll cancel the \(\frac{\pi}{18}\) and get \(h\) by itself Leaaving us with this nice height of \(\large h=1\) :)

OpenStudy (anonymous):

thank you sooo much :)

OpenStudy (luigi0210):

You're welcome ^_^